OpenStudy (anonymous):

Solve the equation by using the quadratic formula:

OpenStudy (anonymous):

\[x^2+4x+3=0\]

OpenStudy (anonymous):

\[x=\frac{ -4 \pm \sqrt{4^2-4*3}}{ 2 }\]\[x=\frac{ -4\pm \sqrt{4} }{ 2 }\]\[x=\frac{ -4\pm2 }{ 2 }\]\[x _{1}=\frac{ -4+2 }{ 2 } = \frac{ -2 }{ 2 } = -1\]\[x _{2}=\frac{ -4-2 }{ 2 }=\frac{ -6 }{ 2 } = -3\]

OpenStudy (anonymous):

so: -1,-3 ??

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

haha(: thanks!! Much appreciated! GBU!! :)